What is the standard cell potential of the following reaction?

2 Ag(s) + I2(s) → 2 I-(aq) + 2 Ag+(aq)

Enter your answer to two decimal places in units of volts.
Answer #1

2 Ag (A) Qage e z ê Ig (A) 12ê – LI (ap) E = -0.60v E= 0.54v 2Ag (A) & Ig (0) - E = – 0.80 +0.54= Agt (as) e QI (ap) -0.26

If you have any questions please comment

If you satisfied with the solution please rate it thanks